OpenStudy (lisa123):

Use mathematical induction to prove the statement is true for all positive integers n, or show why it is false

OpenStudy (lisa123):

4 x 6 + 5 x 7 + 6 x 8 +...+ 4n(4n+2) = 4(4n+1)(8n+7)/6

OpenStudy (lisa123):

I think its false because... 4(4(1)+1)(8(1)+7))/6= 4(5)(15)/6=50

OpenStudy (lisa123):

Pease help @Nnesha

OpenStudy (lisa123):

Please help @dan815

OpenStudy (michele_laino):

if we rewrite your identity with n=1, we get: left side= 4n(4n+2)=4*1(4*1+2)=4*6=24

OpenStudy (michele_laino):

please wait a moment

OpenStudy (lisa123):

I thought that when you did n=1, you were suppose to compare it to the first initial term

OpenStudy (michele_laino):

we have to rewrite the left side setting n=1

OpenStudy (michele_laino):

whereas the right side, with n=1, is: \[\frac{{4\left( {4n + 1} \right)\left( {8n + 1} \right)}}{6} = \frac{{4 \times 5 \times 9}}{6} = 24\]

OpenStudy (lisa123):

Ok what do I do with the 24?

OpenStudy (michele_laino):

since with n=1, the left side is equal to the right side, then our identity is true for n=1

OpenStudy (lisa123):

How is the left side= right side??

OpenStudy (michele_laino):

left side =24 and right side =24 when n=1

OpenStudy (lisa123):

Ohh I think you made a typo because its supposed to be (8n +7) @Michele_Laino

OpenStudy (michele_laino):

it is true, sorry

OpenStudy (michele_laino):

then, the right side is: \[\frac{{4\left( {4n + 1} \right)\left( {8n + 7} \right)}}{6} = \frac{{4 \times 5 \times 15}}{6} = 50\]

OpenStudy (michele_laino):

so, since the left side is 24 whereas the right side is 50 then we can conclude that your identity is false

OpenStudy (lisa123):

Ok makes perfect sense thank you!

OpenStudy (michele_laino):

thank you! :)

OpenStudy (lisa123):

Do you mind helping me with another problem @Michele_Laino ?

OpenStudy (michele_laino):

yes!

OpenStudy (lisa123):

1^2 + 4^2 + 7^2 + ... + (3n - 2)^2 = n(6n^2-3n-1)/2

OpenStudy (michele_laino):

here we have to apply the same procedure of your previous exercise. So I set n=1, and I write your left side:

OpenStudy (lisa123):

Ok (3(1)- 2)^2= 1^2

OpenStudy (michele_laino):

that's right! \[{\left( {3n - 2} \right)^2} = {\left( {3 - 2} \right)^2} = {1^2} = 1\]

OpenStudy (michele_laino):

now I write your right side for n=1:

OpenStudy (michele_laino):

\[\frac{{n\left( {6{n^2} - 3n - 1} \right)}}{2} = \frac{{1 \times \left( {6 - 3 - 1} \right)}}{2} = \frac{{1 \times 2}}{2} = 1\]

OpenStudy (lisa123):

Ok so 1(6(1)^2-3(1)-1)/2= 6-3-1/2 = 2/2=1 yes got same answer

OpenStudy (michele_laino):

ok! So your identity is true for n=1. Now let's suppose your identity true for a generic n

OpenStudy (lisa123):

ok that's where I do n=k. right?

OpenStudy (michele_laino):

yes!

OpenStudy (michele_laino):

Now I write the left side for n+1, namely: \[\begin{gathered} {1^2} + {4^2} + {7^2} + ... + {\left( {3n - 2} \right)^2} + {\left( {3\left( {n + 1} \right) - 2} \right)^2} = \hfill \\ = {1^2} + {4^2} + {7^2} + ... + {\left( {3n - 2} \right)^2} + {\left( {3n + 3 - 2} \right)^2} = \hfill \\ = {1^2} + {4^2} + {7^2} + ... + {\left( {3n - 2} \right)^2} + {\left( {3n + 1} \right)^2} \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

\[\Large \begin{gathered} {1^2} + {4^2} + {7^2} + ... + {\left( {3n - 2} \right)^2} + {\left( {3\left( {n + 1} \right) - 2} \right)^2} = \hfill \\ = {1^2} + {4^2} + {7^2} + ... + {\left( {3n - 2} \right)^2} + {\left( {3n + 3 - 2} \right)^2} = \hfill \\ = {1^2} + {4^2} + {7^2} + ... + {\left( {3n - 2} \right)^2} + {\left( {3n + 1} \right)^2} \hfill \\ \end{gathered} \]

OpenStudy (lisa123):

Ok so for the right side it is (3n-2)^2 +(3n+1)^2 = n+1 (6(n+1)^2-3(n+1)-1)/2

OpenStudy (michele_laino):

that's right! we have: \[\Large \begin{gathered} {1^2} + {4^2} + {7^2} + ... + {\left( {3n - 2} \right)^2} + {\left( {3\left( {n + 1} \right) - 2} \right)^2} = \hfill \\ = {1^2} + {4^2} + {7^2} + ... + {\left( {3n - 2} \right)^2} + {\left( {3n + 3 - 2} \right)^2} = \hfill \\ = {1^2} + {4^2} + {7^2} + ... + {\left( {3n - 2} \right)^2} + {\left( {3n + 1} \right)^2} = \hfill \\ \hfill \\ = \frac{{n\left( {6{n^2} - 3n - 1} \right)}}{2} + {\left( {3n + 1} \right)^2} = \hfill \\ \end{gathered} \]

OpenStudy (lisa123):

Ok this is the part where I get confused as to what's next

OpenStudy (lisa123):

Do I have to simplify further?

OpenStudy (michele_laino):

\[\begin{gathered} \frac{{n\left( {6{n^2} - 3n - 1} \right)}}{2} + {\left( {3n + 1} \right)^2} = \hfill \\ = \frac{{6{n^3} + 15{n^2} + 11n + 2}}{2} = \frac{{\left( {n + 1} \right)\left( {6{{\left( {n + 1} \right)}^2} - 3\left( {n + 1} \right) - 1} \right)}}{2} \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

now, we can note that the last expression, is the right side, of your original identity, with n+1 in place of n

OpenStudy (lisa123):

Oh ok now I see! So I have to prove that the left and right sides are equal. Right?

OpenStudy (michele_laino):

yes! we can state that your original identity is true for all \[n \in \mathbb{N}\]

OpenStudy (lisa123):

What do the symbols above mean?

OpenStudy (michele_laino):

\[n \in \mathbb{N}\] means for n which belong to the set of all natural number, or the set of whole number. That set is indicated, with this symbol: \[\Large \mathbb{N}\]

OpenStudy (michele_laino):

belongs*

OpenStudy (lisa123):

Ok so all of the steps above proved the statement true?

OpenStudy (michele_laino):

yes! generally, the standard procedure, is to verify that your identity for n=2 and for n=3 also

OpenStudy (lisa123):

oh wait can you help me simplify n+1 (6(n+1)^2-3(n+1)-1)/2....im not getting it to equal the left side

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

we can write: \[\large \begin{gathered} \frac{{\left( {n + 1} \right)\left( {6{{\left( {n + 1} \right)}^2} - 3\left( {n + 1} \right) - 1} \right)}}{2} = \hfill \\ \hfill \\ \frac{{\left( {n + 1} \right)\left( {6\left( {{n^2} + 2n + 1} \right) - 3\left( {n + 1} \right) - 1} \right)}}{2} = \hfill \\ \hfill \\ = \frac{{\left( {n + 1} \right)\left( {6{n^2} + 12n + 6 - 3n - 3 - 1} \right)}}{2} = \hfill \\ \hfill \\ = \frac{{\left( {n + 1} \right)\left( {6{n^2} + 9n + 2} \right)}}{2} = \hfill \\ \hfill \\ = \frac{{6{n^3} + 9{n^2} + 2n + 6{n^2} + 9n + 2}}{2} = \hfill \\ \hfill \\ = \frac{{6{n^3} + 15{n^2} + 11n + 2}}{2} = \hfill \\ \end{gathered} \]

OpenStudy (lisa123):

Wow I was distributing it in a weird way to where I got 6k^ 2 +12k+6.....Now I see what you did

OpenStudy (lisa123):

Thank you so much @Michele_Laino for helping me solve the problems step by step!

OpenStudy (michele_laino):

Thank you! :) @lisa123